# A measurement-based protocol for the generation of delocalised quantum states of a mechanical system

### v1

@@ IV.1 Pulsed drive
The conditional mechanical state and the heralding probability in the pulsed scenario can be expressed applying to the output state eq. 15 the POVM LABEL:eq:POVM. The probability of a click can be expressed in the closed form as
@@
where $L_{n}(x)$ are the Laguerre polynomials of $n$-th degree. In Fig. fig. 2 we report the detection probability  eq. 27 and the negativity of the conditional state of the mirror for different pulses durations and initial state thermal occupation number ${n}_{0}$. For short pulse durations, the dynamics produce strongly nonclassical states, though the probability of conditioning on such events is low. In contrast, longer pulses lead to states with Wigner functions that are less negative. This reflects a general trade-off between the degree of nonclassicality and the likelihood of detection. The underlying mechanism is the anti-Stokes photon-phonon generation. Shorter pulses stimulate the emission of only few photons, whose detection results in more nonclassical reduced states. This highlight the fact that more entanglement is no guarantee of more nonclassicality. The negativity of the WF depends on how nonclassical is the detected light, how indicated by fig. 3 c). Fewer photons require more sensible detectors and more runs, possibly yielding nearly single Fock mechanical states. Longer pulses yields modes containing more photons and a higher effective temperature of the mechanical component, thus generating negativity only for efficient pre-cooling.
@@ IV.2 Continuous drive
The detection probability is written in terms of the gaussian WF $W_{\text{opt}}(\zeta)$, depending on $(\bm{\Sigma}_{\text{opt}})^{-1}$, the inverse of the $2x2$ submatrix of $\bm{\Sigma}_{\text{out}}=\begin{pmatrix}\bm{\Sigma}_{m}&\bm{\Sigma}_{c}\\ \bm{\Sigma}_{c}^{T}&\bm{\Sigma}_{\text{opt}}\end{pmatrix}$ relative to output field quadratures $X_{\text{out}}$, $Y_{\text{out}}$. The conditional mechanical WF eq. 32 can be analytically calculated by standard gaussian integrations. The conditional mechanical state
@@
Figure 5: WFs obtained filtering on the Stokes sideband $\nu=-1$, for different modes duration $\Omega\tau=5,10$. The heralding probaiblities are $2.2\%$ and $3.7\%$ respectively. Detection of longer pulses heralds less negative WF due to the higher number of photons. Differently from the pulsed scenario, the length of the detection window is limited sideband resolution, limiting heralding probability. Parameters are the same as in fig. 1.
V Robustness to higher temperatures
V.1 Pulsed regime
$\begin{cases}\dot{\hat{\delta a}}=-\kappa\hat{\delta a}+\frac{ig}{2}\hat{\delta b}^{\dagger}-\sqrt{2\kappa}\hat{a}_{\text{in}},\\ \dot{\hat{\delta b}}=\frac{ig}{2}\hat{\delta a}^{\dagger}-\Gamma\hat{\delta b}-\sqrt{2\kappa}\hat{b}_{\text{in}},\end{cases}$ (35)
where the stochastic thermal input noise $\hat{b}_{\text{in}}(t)$ is defined by $\langle\hat{b}_{\text{in}}(t)\hat{b}^{\dagger}_{\text{in}}(t^{\prime})\rangle=(\bar{n}+1)\delta(t-t^{\prime})$, $\langle\hat{b}^{\dagger}_{\text{in}}(t)\hat{b}_{\text{in}}(t^{\prime})\rangle=\bar{n}\delta(t-t^{\prime})$, and $\bar{n}=(e^{\hbar\Omega/k_{B}T}-1)^{-1}$ [2]. One can then apply a similar treatment to the last section: perform an adiabatic elimination of the cavity allows the definition of the output optical field $\hat{a}_{\text{out}}=\hat{a}_{\text{in}}+\sqrt{2\kappa}~\hat{\delta a}$. By defining appropriate input and output bosonic modes, one finally arrives at
@@ Appendix B Validity of RWA
From this equation it is evident how the resonant term evolves linearly with time with $gt$, while the off-resonant ones oscillates with bounded amplitude by $g/\Omega$. Therefore a first bound on the validity of the RWA can be placed by requiring $gt\gg g/\Omega$, that means $t\gg 1/\Omega$. This implies that if $g/\Omega\ll 1$, after the first few oscillations the off-resonant amplitudes are in magnitude way smaller that the resonant one. However, the non-commutative nature of these two terms imply that for longer times off-resonant effects can build up, invalidating the RWA. This results clear if one approximate the evolution factorizing the resonant and off-resonant interaction
@@ Appendix D Bogoliubov form Input-Output relations derivation
$\int_{s^{\prime}}^{\tau}ds~\psi_{\text{out}}(s)\left[-2\delta(s-s^{\prime})-2Ge^{\mathcal{G}(s-s^{\prime})}\right]=\psi_{\text{in}}(s^{\prime}).$ (57)
, where in the derivation we have used the integration domain property $\int_{0}^{\tau}ds\int_{s}^{\tau}ds^{\prime}=\int_{0}^{\tau}ds^{\prime}\int_{s^{\prime}}^{\tau}ds$. Adopting the ansatz $\psi_{\text{out}}(s)\propto e^{\lambda s}$ yields
@@
$\hat{A}_{\text{in}}=\int_{0}^{\tau}\psi_{\text{in}}(s)\hat{a}_{\text{in}}(s)\hskip 11.38092pt\hat{A}_{\text{out}}=\int_{0}^{\tau}\psi_{\text{out}}(s)\hat{a}_{\text{out}}(s)$ (59)
$\hat{E}_{\text{1,in}}=\int_{0}^{\tau}\psi_{\text{in}}(s)\hat{b}_{\text{in}}(s)\hskip 11.38092pt\hat{E}_{\text{2,in}}=\int_{0}^{\tau}\psi_{\text{out}}(s)\hat{b}_{\text{in}}(s)$

### v2

@@ IV.1 Pulsed drive
The conditional mechanical state and the heralding probability in the pulsed scenario can be expressed by applying to the output state eq. 15 the POVM LABEL:eq:POVM. The probability of a click can be expressed in the closed form as
@@
where $L_{n}(x)$ are the Laguerre polynomials of $n$-th degree. In fig. 3 we report the detection probability  eq. 27 and the negativity of the conditional state of the mirror for different pulses durations and initial state thermal occupation number ${n}_{0}$. For short pulse durations, the dynamics produce strongly nonclassical states, though the probability of conditioning on such events is low. In contrast, longer pulses lead to states with Wigner functions that are less negative. This reflects a general trade-off between the degree of nonclassicality and the likelihood of detection. The underlying mechanism is the anti-Stokes photon-phonon generation. Shorter pulses stimulate the emission of only few photons, whose detection results in more nonclassical reduced states. This highlight the fact that more entanglement is no guarantee of more nonclassicality. The negativity of the WF depends on how nonclassical the detected light is, indicated by fig. 4 c). Fewer photons require more sensible detectors and more runs, possibly yielding nearly single Fock mechanical states. Longer pulses yields modes containing more photons and a higher effective temperature of the mechanical component, thus generating negativity only for efficient pre-cooling.
@@ IV.2 Continuous drive
The detection probability is written in terms of the gaussian WF $W_{\text{opt}}(\zeta)$, depending on $(\bm{\Sigma}_{\text{opt}})^{-1}$, the inverse of the $2\times 2$ submatrix of $\bm{\Sigma}_{\text{out}}=\begin{pmatrix}\bm{\Sigma}_{m}&\bm{\Sigma}_{c}\\ \bm{\Sigma}_{c}^{T}&\bm{\Sigma}_{\text{opt}}\end{pmatrix}$ relative to output field quadratures $X_{\text{out}}$, $Y_{\text{out}}$. The conditional mechanical WF eq. 32 can be analytically calculated by standard gaussian integrations. The conditional mechanical state
@@
Figure 6: WFs obtained filtering on the Stokes sideband $\nu=-1$, for different modes duration $\Omega\tau=5,10$. The heralding probabilities are $2.2\%$ and $3.7\%$ respectively. Detection of longer pulses heralds less negative WF due to the higher number of photons. Differently from the pulsed scenario, the length of the detection window is limited sideband resolution, limiting heralding probability. Parameters are the same as in fig. 2.
@@
V Robustness to higher temperatures
V.1 Pulsed regime
$\begin{cases}\dot{\hat{\delta a}}=-\kappa\hat{\delta a}+\frac{ig}{2}\hat{\delta b}^{\dagger}-\sqrt{2\kappa}\hat{a}_{\text{in}},\\ \dot{\hat{\delta b}}=\frac{ig}{2}\hat{\delta a}^{\dagger}-\Gamma\hat{\delta b}-\sqrt{2\Gamma}\hat{b}_{\text{in}},\end{cases}$ (35)
where the stochastic thermal input noise $\hat{b}_{\text{in}}(t)$ is defined by $\langle\hat{b}_{\text{in}}(t)\hat{b}^{\dagger}_{\text{in}}(t^{\prime})\rangle=(\bar{n}+1)\delta(t-t^{\prime})$, $\langle\hat{b}^{\dagger}_{\text{in}}(t)\hat{b}_{\text{in}}(t^{\prime})\rangle=\bar{n}\delta(t-t^{\prime})$, and $\bar{n}=(e^{\hbar\Omega/k_{B}T}-1)^{-1}$ [2]. One can then apply a similar treatment to the last section: performing an adiabatic elimination of the cavity allows the definition of the output optical field $\hat{a}_{\text{out}}=\hat{a}_{\text{in}}+\sqrt{2\kappa}~\hat{\delta a}$. By defining appropriate input and output bosonic modes, one finally arrives at
@@ Acknowledgements.
Availability of data and materials
The Wolfram Mathematica code can be made available under reasonable request.
Competing interests
The author declares that they have no competing interests.
@@ Appendix B Validity of RWA
From this equation it is evident how the resonant term evolves linearly with time with $gt$, while the off-resonant ones oscillates with bounded amplitude by $g/\Omega$. Therefore a first bound on the validity of the RWA can be placed by requiring $gt\gg g/\Omega$, that means $t\gg 1/\Omega$. This implies that if $g/\Omega\ll 1$, after the first few oscillations the off-resonant amplitudes are significantly smaller in magnitude than the resonant oscillations. However, the non-commutative nature of these two terms imply that for longer times off-resonant effects can build up, invalidating the RWA. This is clear if one approximate the evolution factorizing the resonant and off-resonant interaction
@@ Appendix D Bogoliubov form Input-Output relations derivation
$\int_{s^{\prime}}^{\tau}ds~\psi_{\text{out}}(s)\left[-2\delta(s-s^{\prime})-2Ge^{\mathcal{G}(s-s^{\prime})}\right]=\psi_{\text{in}}(s^{\prime}),$ (57)
where in the derivation we have used the integration domain property $\int_{0}^{\tau}ds\int_{s}^{\tau}ds^{\prime}=\int_{0}^{\tau}ds^{\prime}\int_{s^{\prime}}^{\tau}ds$. Adopting the ansatz $\psi_{\text{out}}(s)\propto e^{\lambda s}$ yields
@@
$\begin{aligned} &\hat{A}_{\text{in}}=\int_{0}^{\tau}\psi_{\text{in}}(s)\hat{a}_{\text{in}}(s)\hskip 11.38092pt\hat{A}_{\text{out}}=\int_{0}^{\tau}\psi_{\text{out}}(s)\hat{a}_{\text{out}}(s)\\ &\hat{E}_{\text{1,in}}=\int_{0}^{\tau}\psi_{\text{in}}(s)\hat{b}_{\text{in}}(s)\hskip 11.38092pt\hat{E}_{\text{2,in}}=\int_{0}^{\tau}\psi_{\text{out}}(s)\hat{b}_{\text{in}}(s)\end{aligned}.$ (59)